# TsCA: On the Semantic Consistency Alignment via Conditional Transport for Compositional Zero-Shot Learning

### v1

@@ Abstract
Compositional Zero-Shot Learning (CZSL) aims to recognize novel state-object compositions by leveraging the shared knowledge of their primitive components. Despite considerable progress, effectively calibrating the bias between semantically similar multimodal representations, as well as generalizing pre-trained knowledge to novel compositional contexts, remains an enduring challenge. In this paper, our interest is to revisit the conditional transport (CT) theory and its homology to the visual-semantics interaction in CZSL and further, propose a novel Trisets Consistency Alignment framework (dubbed TsCA) that well-addresses these issues. Concretely, we utilize three distinct yet semantically homologous sets, i.e., patches, primitives, and compositions, to construct pairwise CT costs to minimize their semantic discrepancies. To further ensure the consistency transfer within these sets, we implement a cycle-consistency constraint that refines the learning by guaranteeing the feature consistency of the self-mapping during transport flow, regardless of modality. Moreover, we extend the CT plans to an open-world setting, which enables the model to effectively filter out unfeasible pairs, thereby speeding up the inference as well as increasing the accuracy. Extensive experiments are conducted to verify the effectiveness of the proposed method.
Introduction
Identifying new concepts from a set of seen primitives is one of the fundamental challenges for AI systems to imitate the human learning process. Imagine a new cuisine—Spicy Chocolate Cake. Although it may sound like an odd combination, one can still perceive its appearance and taste based on existing knowledge of ‘spiciness’ and ‘Chocolate’. Likewise, compositional zero-shot learning (CZSL)  (Misra, Gupta, and Hebert 2017; Naeem et al. 2021) emerges as a compelling paradigm that seeks to endow machines with similar cognitive abilities, allowing them to understand the composed primitives (i.e., states and objects) during training and generalize to unseen compositions for inference.
Technically, studies on CZSL have revolved around achieving fine-grained alignment across the image and composed state-object text domains (Huynh and Elhamifar 2020). Pioneer methods typically load pre-trained image encoders and word embeddings to extract multi-modal features. Various alignment tricks are then applied to constrain the shared latent space, including concept learning (Xu, Kordjamshidi, and Chai 2021), geometric properties (Li et al. 2020), semantic transformation (Nagarajan and Grauman 2018), and graph embedding (Mancini et al. 2022). Moving beyond these post-processing alignments, a series of prompt-tuning methods have been developed to explore pre-trained vision-language models (VLMs) for the CZSL scenario (Xu et al. 2022; Nayak, Yu, and Bach 2023), achieving state-of-the-art prediction performance. Pre-trained on massive amounts of image-text pairs, VLMs (e.g., CLIP (Radford et al. 2021)) show impressive abilities to match the visual image with its textual label (prompt) in the multimodal space. Those models often employ a multi-path paradigm for fine-grained textual features, e.g., one composition and two primitive branches. One of the core challenges ensues: How to align the image with such multiple textual labels? Recent studies attempt to address it from different fields, such as hierarchical prompt searching (Wang et al. 2023a), decomposed feature fusion (Lu et al. 2023a), and multi-step observation (Li et al. 2023a). Although showing attractive results, these methods focus primarily on image-to-composition (or to-primitive) alignments, ignoring intrinsic relationships within the composition and its primitives. This may fail to capture semantic consistency among the image-composition-primitive interactions, leading to suboptimal alignments.
@@ Introduction
To address the above issues, this paper proposes the Trisets Consistency Alignment framework (dubbed TsCA), which is built upon a novel Consistency-aware Conditional Transport (CCT) theory derived from the new view of CZSL. As shown in Fig. 1, we represent an image in three directions, i.e., ${\bf P}_{1}$, a distribution over all visual patches; ${\bf P}_{2}$, a distribution over the composition set; and ${\bf P}_{3}$, a distribution over the primitive set. Concretely, ${\bf P}_{1}$ captures the detailed visual features of an image, while ${\bf P}_{2}$ and ${\bf P}_{3}$ denote the global and local textual concepts of the same content. Therefore, the task of CZSL can be viewed as aligning these three discrete distributions as closely as possible. Accordingly, it is indeed key to properly measure the distance between empirical distributions with different supports. Fortunately, conditional transport (CT), well-examined in recent research, offers a bidirectional measurement of the distance from one distribution to another, given the cost matrix. Inheritedly, TsCA extends the minimization of CT to the alignment of our cross-modal distribution trio.
@@
Specifically, TsCA gracefully facilitates the matching of image-composition-primitive triplets by meticulously crafting three pairs of CTs, thereby aligning the rich tapestry of cross-modal and intra-modal semantics. First, $\mathcal{CT}({\bf P}_{1},{\bf P}_{2})$ measures the transport distance between the visual path set ${\bf P}_{1}$ and the composition set ${\bf P}_{2}$. On the one hand, it helps the label be transported to the compositional visuals with higher probabilities, and on the other hand, it regularizes the finetuning for better cross-modal alignments. Next, $\mathcal{CT}({\bf P}_{1},{\bf P}_{3})$ focuses on discovering the corresponding attribute and object patches from the disentangle perspective. Last, unlike the above two, $\mathcal{CT}({\bf P}_{2},{\bf P}_{3})$ aims at intra-modal interactions, which are often overlooked by previous studies. By minimizing the transport distance between the composition and its primitives explicitly, the textual outputs are assumed to present high semantic coherence. E.g., the composition vectors are closer to their primitives in the embedding space. Moreover, the learned transport plan in $\mathcal{CT}({\bf P}_{2},{\bf P}_{3})$ further provides an efficient tool to filter out unfeasible compositions, which benefits the open-world setting.
@@
More importantly, as discussed above, TsCA formulates the CZSL as the alignment of three distributions and seeks to minimize their transport costs. An intuitive solution is to run pairwise transportation and optimize the above three CT costs. Unfortunately, this case could not model the complex relations among these sets. To capture deeper interactions, we extend the naive CT to a well-designed consistency-aware CT (CCT) for the CZSL triplet case. Motivated by the fact that these three distributions describe the same semantics, CCT regularizes the product of three transport plans as a diagonal matrix from two directions (shown in Fig. 1). In other words, the composition label will return to itself after a cycle of transportation, which ensures semantic consistency across sets.
@@
Related Work
Compositional Zero-Shot Learning
CZSL(Mancini et al. 2021) is a specialized case of zero-shot learning (ZSL) (Liu et al. 2023c). Given the same set of objects with associated states, it focuses on recognizing unseen state-object compositions by learning from seen compositions. In general, traditional CZSL can be divided into two streams, i.e., compositional classification and simple primitive classification. The former directly predicts compositions by aligning visual features and composed labels in a shared space, or resorting to a graph network for contextuality modeling (Tanwisuth et al. 2023; Saini, Pham, and Shrivastava 2022; Li et al. 2022). Conversely, the latter identifies states and objects independently and constructs the joint compositional probability distribution. They either disregard the contextuality between primitives or impose training-specific correlations that are detrimental to generalization.
Recently, equipped with prompt learning, large-scale VLMs like CLIP are empowered to adapt to CZSL by reducing domain shift and leveraging pre-trained knowledge. Further, researchers attempt to explore merging those two typical paradigms to create an integrated multi-path paradigm. For instance, HPL (Wang et al. 2023a) learns three hierarchical prompts by explicitly fixing the unrelated word tokens in the three embedding spaces at different levels. Troika (Huang et al. 2024) effectively aligns the branch-specific prompt representations and decomposed visual features with a cross-modal traction module. Our TsCA aligns closely with this paradigm, although with a greater emphasis on direct inquiries into semantic alignment within distributions.
Conditional Transport
Recently, CT has acted as an efficient tool to measure the distance between two discrete distributions (Zheng and Zhou 2021). Through learning bidirectional transport plans based on semantic similarity between samples from both source and target distributions, CT achieves fine-grained matching with the mini-batch optimization. More importantly, CT can integrate seamlessly with deep-learning frameworks and its effectiveness has been widely demonstrated in various applications (Li et al. 2023b; Wang et al. 2022a; Tanwisuth et al. 2023). For example, (Tian et al. 2023) propose a novel CT-based imbalanced transductive few-shot learning model to fully exploit unbiased statistics of imbalanced query samples. Similarly, (Tanwisuth et al. 2021) employs a probabilistic bi-directional transport between target features and class prototypes for unsupervised domain adaptation, showcasing robustness against class imbalance and facilitating domain adaptation without direct access to the source data. For the first time, through empirical demonstration, we establish that CT, boasting attributes like lower complexity and better scalability, is equally viable for CZSL.
@@
Problem Formulation
Given state set $\mathcal{S}=\{s_{0},s_{1},...,s_{|\mathcal{S}|}\}$ together with object set $\mathcal{O}=\{o_{0},o_{1},...,o_{|\mathcal{O}|}\}$ , we can define the label space with their Cartesian product, $\mathcal{C}=\mathcal{S}\times\mathcal{O}$. Then $\mathcal{C}$ can be divided into two disjoint label subsets such that $\mathcal{C}^{se}\in\mathcal{C}$, $\mathcal{C}^{us}\in\mathcal{C}$, and $\mathcal{C}^{se}\cap\mathcal{C}^{us}=\emptyset$ where $\mathcal{C}^{se}$ and $\mathcal{C}^{us}$ are the set of the seen and unseen classes respectively. Specifically, during the training phase, $\mathcal{C}^{se}$ are used to train a discriminative model $\mathcal{M}:\mathcal{X}\overrightarrow{\mathcal{missing}}{C}^{se}$ from the input image space to candidate composition label set. At inference time, the model is expected to predict unseen compositions in the test sample space, i.e.,$\mathcal{M}:\mathcal{X}\overrightarrow{\mathcal{missing}}{C}^{test}$. In this paper, we follow the setting of Generalized ZSL (Xian et al. 2017), considering testing samples contain both seen and unseen compositions. In general, in the closed-world evaluation, only the known composition space of test samples is required as $\mathcal{C}^{test}=\mathcal{C}^{se}\cup\mathcal{C}^{us}$. For the open-world evaluation (Karthik, Mancini, and Akata 2022), the model has to consider all possible permutations of the state-object pairs, i.e., $\mathcal{C}^{test}=\mathcal{C}$.
Image as Three Sets
Built upon the pre-trained CLIP and with the soft prompt tuning technique, our TsCA represents an image as three sets: visual patch set ${\bf P}_{1}$, textual composition set ${\bf P}_{2}$, and primitive set ${\bf P}_{3}$. These sets capture the multimodal features of the same content from different semantic domains, acting as a fundamental role in the fine-grained alignments.
@@
For an input image $x$, CLIP first splits it into $N$ non-overlapping patches evenly and then feeds them (with a [CLS] token inserted) into the image encoder to extract the embedding of the [CLS] token $\boldsymbol{x}^{c}_{CLS}$ and the patch features $\boldsymbol{x}=\{\boldsymbol{x}_{n}|_{n=1}^{N}\}\in\mathbb{R}^{d\times N}$, where $\boldsymbol{x}_{n}\in\mathbb{R}^{d}$ denotes the embedding of patch $n$ with the embedding dimension being $d$. Naturally, the discrete probability distribution over the patch set can be formulated as:
${\bf P}_{1}=\sum_{n=1}^{N}\theta_{n}\delta_{\boldsymbol{x}_{n}},\quad\theta_{n}=\frac{1}{N},$ (1)
where $\delta$ denotes the Dirac delta function. We view all the patches equally and employ the Uniform distribution to model the patch weight $\boldsymbol{\theta}$. Note that ${\bf P}_{1}$ collects the detailed visual features of the local region, thereby bringing benefits to discriminative representation learning, especially when different concepts require emphasis on distinct areas within the input images. Besides, the [CLS] visual token $\boldsymbol{x}_{CLS}^{c}$ is also learned, serving as the global representation, which is often used as the image feature to downstream tasks (Zhou et al. 2022a). Here, we consider it as the visual prior that constructs subsequent textual sets.
@@
In the textual domain, we build composition labels into a learnable soft prompt [v1][v2][v3][state][object], where [v] is the prefix vectors, and [state][object] denote the state and object name, respectively. Then, the prompt will be fed into the text encoder to obtain textual representations $\boldsymbol{y}=\{\boldsymbol{y}_{m}|_{m=1}^{M}\}\in\mathbb{R}^{d\times M}$, where $M$ is the number of training compositions. To embrace the variety of content and minimize the cross-modal disparities, we follow previous work (Huang et al. 2024) and incorporate a residual component obtained through a cross-attention mechanism (Vaswani 2017) with patch embeddings (we here still use $\boldsymbol{y}$ as the output compositions):
$\boldsymbol{y}=\text{Cross-Att}(\boldsymbol{y}^{in},\boldsymbol{x},\boldsymbol{x})+\boldsymbol{y}^{in},$ (2)
where $\boldsymbol{y}^{in}$ denotes the input compositions. Cross-Att is the cross-attention layer with the query $\boldsymbol{y}^{in}$, key $\boldsymbol{x}$, and value $\boldsymbol{x}$ as inputs, and outputs the fused features. As a result, the composition set can be viewed as:
${\bf P}_{2}=\sum_{m=1}^{M}\alpha_{m}\delta_{\boldsymbol{y}_{m}},\quad\boldsymbol{\alpha}=\sigma(\boldsymbol{y}^{T}\boldsymbol{x}^{c}_{CLS}),$ (3)
where $\sigma$ denotes the softmax function. We calculate the composition weights $\boldsymbol{\alpha}$ via the semantic similarity of the composition label and visual feature. This helps ${\bf P}_{2}$ to focus on compositions that describe the input image well.
@@
Unlike the composition set that focuses on global textual features, the primitive set aims to explore the disentangled state and object representations. Motivated by the recent multi-path paradigms, we learn the corresponding primitive representations through [v1][v2][v3][state] and [v1][v2][v3][object], with the outputs denoted as $\boldsymbol{z}^{s}\in\mathbb{R}^{d\times{|\mathcal{S}|}}$ and $\boldsymbol{z}^{o}\in\mathbb{R}^{d\times{|\mathcal{O}|}}$ respectively. The corresponding probability weights are calculated as:
$\boldsymbol{\beta}^{s}=\sigma(({\boldsymbol{z}^{s}})^{T}\boldsymbol{x}^{s}_{CLS}),\quad\boldsymbol{\beta}^{o}=\sigma(({\boldsymbol{z}^{o}})^{T}\boldsymbol{x}^{o}_{CLS}),$ (4)
where $\boldsymbol{x}^{s}_{CLS}$ and $\boldsymbol{x}^{o}_{CLS}$ are two adapted visual features, which can be obtained via a lightweight adapter $g$:
$[\boldsymbol{x}^{s}_{CLS},\boldsymbol{x}^{o}_{CLS}]=g(\boldsymbol{x}^{c}_{CLS}),$ (5)
where $g$ takes the image feature $\boldsymbol{x}^{c}_{CLS}$ as input, and outputs the state/object-relevant features, deriving unique visual representations. To complete the primitive set, we concatenate the feature points from both state and object labels as $\boldsymbol{z}=\{\boldsymbol{z}_{k}|_{k=1}^{K}\}\in\mathbb{R}^{d\times K}$, where $K=|\mathcal{S}|+|\mathcal{O}|$ is the total number of state and object labels. Like compositions that fuse the patch features with Eq. 2, $\boldsymbol{z}$ is fed into the same cross-attention layer:
$\boldsymbol{z}=\text{Cross-Att}(\boldsymbol{z}^{in},\boldsymbol{x},\boldsymbol{x})+\boldsymbol{z}^{in}.$ (6)
@@
${\bf P}_{3}=\sum_{k=1}^{K}\beta_{k}\delta_{\boldsymbol{z}_{k}},\quad\boldsymbol{\beta}=\sigma(\boldsymbol{\beta}^{s}\oplus\boldsymbol{\beta}^{o}),$ (7)
@@
Semantic Consistency Alignment

### v3

@@ Abstract
Compositional Zero-Shot Learning (CZSL) aims to recognize novel state-object compositions by leveraging the shared knowledge of their primitive components. Despite considerable progress, effectively calibrating the bias between semantically similar multimodal representations, as well as generalizing pre-trained knowledge to novel compositional contexts, remains an enduring challenge. In this paper, our interest is to revisit the conditional transport (CT) theory and its homology to the visual-semantics interaction in CZSL and further, propose a novel Trisets Consistency Alignment framework (dubbed TsCA) that well-addresses these issues. Concretely, we utilize three distinct yet semantically homologous sets, i.e., patches, primitives, and compositions, to construct pairwise CT costs to minimize their semantic discrepancies. To further ensure the consistency transfer within these sets, we implement a cycle-consistency constraint that refines the learning by guaranteeing the feature consistency of the self-mapping during transport flow, regardless of modality. Moreover, we extend the CT plans to an open-world setting, which enables the model to effectively filter out unfeasible pairs, thereby speeding up the inference as well as increasing the accuracy. Extensive experiments are conducted to verify the effectiveness of the proposed method [^1] Code is available at: https://anonymous.4open.science/r/TsCA-0E7A..
1 Introduction
Identifying new concepts from a set of seen primitives is one of the fundamental challenges for AI systems to imitate the human learning process. Imagine a new cuisine—Spicy Chocolate Cake. Although it may sound like an odd combination, one can still perceive its appearance and taste based on existing knowledge of ‘spiciness’ and ‘Chocolate’. Likewise, compositional zero-shot learning (CZSL)  Misra et al. (2017); Naeem et al. (2021) emerges as a compelling paradigm that seeks to endow machines with similar cognitive abilities, allowing them to understand the composed primitives (i.e., states and objects) during training and generalize to unseen compositions for inference.
@@ 1 Introduction
Technically, studies on CZSL have revolved around achieving fine-grained alignment across the image and composed state-object text domains Huynh and Elhamifar (2020). Pioneer methods typically load pre-trained image encoders and word embeddings to extract multi-modal features. Various alignment tricks are then applied to constrain the shared latent space, including concept learning Xu et al. (2021), geometric properties Li et al. (2020), semantic transformation Nagarajan and Grauman (2018), and graph embedding Mancini et al. (2022). Moving beyond these post-processing alignments, a series of prompt-tuning methods have been developed to explore pre-trained vision-language models (VLMs) for the CZSL scenario Xu et al. (2022); Nayak et al. (2023), achieving state-of-the-art prediction performance. Pre-trained on massive amounts of image-text pairs, VLMs (e.g., CLIP Radford et al. (2021)) show impressive abilities to match the visual image with its textual label (prompt) in the multimodal space. Those models often employ a multi-path paradigm for fine-grained textual features, e.g., one composition and two primitive branches. One of the core challenges ensues: How to align the image with such multiple textual labels? Recent studies attempt to address it from different fields, such as hierarchical prompt searching Wang et al. (2023), decomposed feature fusion Lu et al. (2023), and multi-step observation Li et al. (2023a). Although showing attractive results, these methods focus primarily on image-to-composition (or to-primitive) alignments, ignoring intrinsic relationships within the composition and its primitives. This may fail to capture semantic consistency among the image-composition-primitive interactions, leading to suboptimal alignments.
@@
To address the above issues, this paper proposes the Trisets Consistency Alignment framework (dubbed TsCA), which is built upon a novel Consistency-aware Conditional Transport (CCT) derived from the new view of CZSL. As shown in Fig. 1, we represent an image in three directions, i.e., ${\bf P}_{1}$, a distribution over all visual patches; ${\bf P}_{2}$, a distribution over the composition set; and ${\bf P}_{3}$, a distribution over the primitive set. Concretely, ${\bf P}_{1}$ captures the detailed visual features of an image, while ${\bf P}_{2}$ and ${\bf P}_{3}$ denote the global and local textual concepts of the same content. Therefore, the task of CZSL can be viewed as aligning these three discrete distributions as closely as possible. Accordingly, it is indeed key to properly measure the distance between empirical distributions with different supports. Fortunately, conditional transport (CT), well-examined in recent research, offers a bidirectional measurement of the distance from one distribution to another, given the cost matrix. Inheritedly, TsCA extends the minimization of CT to the alignment of our cross-modal distribution trio.
@@
Specifically, TsCA gracefully facilitates the matching of image-composition-primitive triplets by meticulously crafting three pairs of CTs, thereby aligning the rich tapestry of cross-modal and intra-modal semantics. First, $\mathcal{CT}({\bf P}_{1},{\bf P}_{2})$ measures the transport distance between the visual patch set and the composition set. On the one hand, it helps the label be transported to the compositional visuals with higher probabilities, and on the other hand, it regularizes the finetuning for better cross-modal alignments. Next, $\mathcal{CT}({\bf P}_{1},{\bf P}_{3})$ focuses on discovering the corresponding attribute and object patches from the disentangle perspective. Last, unlike the above two, $\mathcal{CT}({\bf P}_{2},{\bf P}_{3})$ aims at intra-modal interactions, which are often overlooked by previous studies. By minimizing the transport distance between the composition and its primitives explicitly, the textual outputs are assumed to present high semantic coherence. E.g., the composition vectors are closer to their primitives in the embedding space. Moreover, the learned transport plan in $\mathcal{CT}({\bf P}_{2},{\bf P}_{3})$ further provides an efficient tool to filter out unfeasible compositions, which benefits the open-world setting.
@@
More importantly, as discussed above, TsCA formulates the CZSL as the alignment of three distributions and seeks to minimize their transport costs. An intuitive solution is to run pairwise transportation and optimize the above three CT costs. Unfortunately, this case could not model the complex relations among these sets. To capture deeper interactions, we extend the naive CT to a well-designed consistency-aware CT (CCT) for the CZSL triplet case. Motivated by the fact that these three distributions describe the same semantics, CCT regularizes the product of three transport plans as a diagonal matrix from clockwise direction (shown in Fig. 1). In other words, the composition label will return to itself after a cycle of transport, which ensures semantic consistency across sets.
@@
2 Related Work
2.1 Compositional Zero-Shot Learning
CZSLMancini et al. (2021) is a specialized case of zero-shot learning (ZSL) Liu et al. (2023). Given the same set of objects with associated states, it focuses on recognizing unseen state-object compositions by learning from seen compositions. In general, traditional CZSL can be divided into two streams, i.e., compositional classification and simple primitive classification. The former directly predicts compositions by aligning visual features and composed labels in a shared space, or resorting to a graph network for contextuality modeling Tanwisuth et al. (2023); Saini et al. (2022); Li et al. (2022). Conversely, the latter identifies states and objects independently and constructs the joint compositional probability distribution. They either disregard the contextuality between primitives or impose training-specific correlations that are detrimental to generalization.
Recently, equipped with prompt learning, large-scale VLMs like CLIP are empowered to adapt to CZSL by reducing domain shift and leveraging pre-trained knowledge . Further, researchers attempt to explore merging those two typical paradigms to create an integrated multi-path paradigm. For instance, HPL Wang et al. (2023) learns three hierarchical prompts by explicitly fixing the unrelated word tokens in the three embedding spaces at different levels. Troika Huang et al. (2024) effectively aligns the branch-specific prompt representations and decomposed visual features with a cross-modal traction module. Our TsCA aligns closely with this paradigm, although with a greater emphasis on direct inquiries into semantic alignment within distributions.
2.2 Conditional Transport
Recently, CT has acted as an efficient tool to measure the distance between two discrete distributions Zheng and Zhou (2021). Through learning bidirectional transport plans based on semantic similarity between samples from both source and target distributions, CT achieves fine-grained matching with the mini-batch optimization. More importantly, CT can integrate seamlessly with deep-learning frameworks and its effectiveness has been widely demonstrated in various applications Li et al. (2023b); Wang et al. (2022); Tanwisuth et al. (2023). For example, Tian et al. (2023) propose a novel CT-based imbalanced transductive few-shot learning model to fully exploit unbiased statistics of imbalanced query samples. Similarly, Tanwisuth et al. (2021) employs a probabilistic bidirectional transport between target features and class prototypes for unsupervised domain adaptation, showcasing robustness against class imbalance and facilitating domain adaptation without direct access to the source data. For the first time, through empirical demonstration, we establish that CT, boasting attributes like lower complexity and better scalability, is equally viable for CZSL.
@@ 3 Methodology
3.1 Problem Formulation
Given state set $\mathcal{S}=\{s_{0},s_{1},...,s_{|\mathcal{S}|}\}$ together with object set $\mathcal{O}=\{o_{0},o_{1},...,o_{|\mathcal{O}|}\}$ , we can define the label space with their Cartesian product, $\mathcal{C}=\mathcal{S}\times\mathcal{O}$. Then $\mathcal{C}$ can be divided into two disjoint label subsets such that $\mathcal{C}^{se}\in\mathcal{C}$, $\mathcal{C}^{us}\in\mathcal{C}$, and $\mathcal{C}^{se}\cap\mathcal{C}^{us}=\emptyset$ where $\mathcal{C}^{se}$ and $\mathcal{C}^{us}$ are the set of the seen and unseen classes respectively. Specifically, during the training phase, $\mathcal{C}^{se}$ are used to train a discriminative model $\mathcal{M}:\mathcal{X}\overrightarrow{\mathcal{missing}}{C}^{se}$ from the input image space to candidate composition label set. At inference time, the model is expected to predict unseen compositions in the test sample space, i.e.,$\mathcal{M}:\mathcal{X}\overrightarrow{\mathcal{missing}}{C}^{test}$. In this paper, we follow the setting of Generalized ZSL Xian et al. (2017), considering testing samples contain both seen and unseen compositions. In general, in the closed-world evaluation, only the known composition space of test samples is required as $\mathcal{C}^{test}=\mathcal{C}^{se}\cup\mathcal{C}^{us}$. For the open-world evaluation Karthik et al. (2022), the model has to consider all possible permutations of the state-object pairs, i.e., $\mathcal{C}^{test}=\mathcal{C}$.
3.2 Image as Three Sets
Built upon the pre-trained CLIP and with the soft prompt tuning technique, our TsCA represents an image as three sets: visual patch set , textual composition set , and primitive set . These sets capture the multimodal features of the same content from different semantic domains, acting as a fundamental role in the fine-grained alignments.
@@ 3.2 Image as Three Sets
For an input image $x$, CLIP first splits it into $N$ non-overlapping patches evenly and then feeds them (with a [CLS] token inserted) into the image encoder to extract the embedding of the [CLS] token $\bm{x}^{c}_{CLS}$ and the patch features $\bm{x}=\{\bm{x}_{n}|_{n=1}^{N}\}\in\mathbb{R}^{d\times N}$, where $\bm{x}_{n}\in\mathbb{R}^{d}$ denotes the embedding of patch $n$ with the embedding dimension being $d$. Naturally, the discrete probability distribution over the patch set can be formulated as:
${\bf P}_{1}=\sum_{n=1}^{N}\theta_{n}\delta_{\bm{x}_{n}},\quad\theta_{n}=\frac{1}{N},$ (1)
where $\delta$ denotes the Dirac delta function. We view all the patches equally and employ the Uniform distribution to model the patch weight $\bm{\theta}$. Note that ${\bf P}_{1}$ collects the detailed visual features of the local region, thereby bringing benefits to discriminative representation learning, especially when different concepts require emphasis on distinct areas within the input images. Besides, the [CLS] visual token $\bm{x}_{CLS}^{c}$ is also learned, serving as the global representation, which is often used as the image feature to downstream tasks Zhou et al. (2022). Here, we consider it as the visual prior that constructs subsequent textual sets.
@@
In the textual domain, we build composition labels into a learnable soft prompt [v1][v2][v3][state][object], where [v] is the prefix vectors, and [state][object] denote the state and object name, respectively. Then, the prompt will be fed into the text encoder to obtain textual representations $\bm{y}=\{\bm{y}_{m}|_{m=1}^{M}\}\in\mathbb{R}^{d\times M}$, where $M$ is the number of training compositions. To embrace the variety of content and minimize the cross-modal disparities, we follow previous work Huang et al. (2024) and incorporate a residual component obtained through a cross-attention mechanism Vaswani (2017) with patch embeddings (we here still use $\bm{y}$ as the output compositions):
$\bm{y}=\text{Cross-Att}(\bm{y}^{in},\bm{x},\bm{x})+\bm{y}^{in},$ (2)
where $\bm{y}^{in}$ denotes the input compositions. Cross-Att is the cross-attention layer with the query $\bm{y}^{in}$, key $\bm{x}$, and value $\bm{x}$ as inputs, and outputs the fused features. As a result, the composition set can be viewed as:
${\bf P}_{2}=\sum_{m=1}^{M}\alpha_{m}\delta_{\bm{y}_{m}},\quad\bm{\alpha}=\sigma(\bm{y}^{T}\bm{x}^{c}_{CLS}),$ (3)
where $\sigma$ denotes the softmax function. We calculate the composition weights $\bm{\alpha}$ via the semantic similarity of the composition label and visual feature. This helps ${\bf P}_{2}$ to focus on compositions that describe the input image well.
@@
Unlike the composition set that focuses on global textual features, the primitive set aims to explore the disentangled state and object representations. Motivated by the recent multi-path paradigms, we learn the corresponding primitive representations through [v1][v2][v3][state] and [v1][v2][v3][object], with the outputs denoted as $\bm{z}^{s}\in\mathbb{R}^{d\times{|\mathcal{S}|}}$ and $\bm{z}^{o}\in\mathbb{R}^{d\times{|\mathcal{O}|}}$ respectively. The corresponding probability weights are calculated as:
$\bm{\beta}^{s}=\sigma(({\bm{z}^{s}})^{T}\bm{x}^{s}_{CLS}),\quad\bm{\beta}^{o}=\sigma(({\bm{z}^{o}})^{T}\bm{x}^{o}_{CLS}),$ (4)
where $\bm{x}^{s}_{CLS}$ and $\bm{x}^{o}_{CLS}$ are two adapted visual features, which can be obtained via a lightweight adapter $g$:
$[\bm{x}^{s}_{CLS},\bm{x}^{o}_{CLS}]=g(\bm{x}^{c}_{CLS}),$ (5)
where $g$ takes the image feature $\bm{x}^{c}_{CLS}$ as input, and outputs the state/object-relevant features, deriving unique visual representations. To complete the primitive set, we concatenate the feature points from both state and object labels as $\bm{z}=\{\bm{z}_{k}|_{k=1}^{K}\}\in\mathbb{R}^{d\times K}$, where $K=|\mathcal{S}|+|\mathcal{O}|$ is the total number of state and object labels. Like compositions that fuse the patch features with Eq. 2, $\bm{z}$ is fed into the same cross-attention layer:
$\bm{z}=\text{Cross-Att}(\bm{z}^{in},\bm{x},\bm{x})+\bm{z}^{in}.$ (6)
@@
${\bf P}_{3}=\sum_{k=1}^{K}\beta_{k}\delta_{\bm{z}_{k}},\quad\bm{\beta}=\sigma(\bm{\beta}^{s}\oplus\bm{\beta}^{o}),$ (7)
@@
3.3 Semantic Consistency Alignment